OpenStudy (anonymous):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many women auditioned? 44 36 76 60

OpenStudy (anonymous):

is it 36?

OpenStudy (unklerhaukus):

\[m+w=100\]\[3\times w-2\times m=20\]

OpenStudy (anonymous):

so is it 36? im checking myself?

OpenStudy (anonymous):

20 is not an option

OpenStudy (unklerhaukus):

show working?

OpenStudy (unklerhaukus):

that is not what i get

OpenStudy (unklerhaukus):

which method do you want/

OpenStudy (anonymous):

im confused now because 20 is not even an option

OpenStudy (unklerhaukus):

who said 20?

OpenStudy (anonymous):

u did m=20

OpenStudy (unklerhaukus):

wait did you get the same second equation as me?

OpenStudy (anonymous):

3 * w - 2 * m = 20

OpenStudy (anonymous):

i got 36

OpenStudy (unklerhaukus):

m is way more than 20

OpenStudy (unklerhaukus):

in that case i dont know how you got 36, if i use -3 * w + 2 * m = 20 i get 36

OpenStudy (anonymous):

i got 36 too

OpenStudy (unklerhaukus):

but i dont think that is the question

OpenStudy (unklerhaukus):

show your working please

OpenStudy (anonymous):

its timed i gotta go ... thanks

OpenStudy (unklerhaukus):

nevermind